The surface pressure on the planet venus is 95 atm,how far below th surface of the ocean in earth do you need to be to experience the same pressure


1
Expert's answer
2021-02-19T18:57:52-0500
"P=P_{atm}+\rho gd,""P-P_{atm}=\rho gd,""d=\dfrac{P-P_{atm}}{\rho g},""d=\dfrac{(95\ atm-1\ atm)\cdot\dfrac{1.013\cdot10^5\ Pa}{1\ atm}}{1025\ \dfrac{kg}{m^3}\cdot9.8\ \dfrac{m}{s^2}}=948\ m."